A certain test preparation course is designed to help students improve their scores on the GRE exam. A mock exam is given at the beginning and end of the course to determine the effectiveness of the course. The following measurements are the net change in 55 students' scores on the exam after completing the course:

16,12,11,13,26

Using these data, construct a 95% confidence interval for the average net change in a student's score after completing the course. Assume the population is approximately normal.

1.Calculate the sample mean for the given sample data. Round your answer to one decimal place.

2.Calculate the sample standard deviation for the given sample data. Round your answer to one decimal place.

3.Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places.

4.Construct the 95%95% confidence interval. Round your answer to one decimal place.
Answer #1

1) \bar x = (16 + 12 + 11 + 13 + 26)/5 = 15.6

2) s = sqrt(((16 - 15.6)^2 + (12 - 15.6)^2 + (11 - 15.6)^2 + (13 - 15.6)^2 + (26 - 15.6)^2)/4) = 6.1

3) At 95% confidence level, the critical value is t* = 2.777

4) The 95% confidence interval is

\bar x +/- t* * s/п

= 15.6 +/- 2.777 * 6.1/sqrt(5)

= 15.6 +/- 7.6

= 8.0, 23.2